OpenStudy (anonymous):

L'HOPITAL rule Lim x-> Infinite x sin (10/x) Please explain

OpenStudy (anonymous):

Ok: Indeterminate of type 0*infinity So we let t = 1/x then now we have \[\lim_{x \rightarrow infinity} xsin(10/x) = \lim_{t \rightarrow 0} \sin(10t)/(t) \] which is: \[\lim_{t \rightarrow 0} 10\cos(10t) \] from applying L'Hopital. Now plug in t and see that the limit goes to 10.

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

yes it does. thank you!